please do number 3 only, thank you










In Exercises 2-9 find a matrix P that orthogonally diagonalizes A, and determine p-AP 3. A = 213 ]
Answer #1

Sel: BS Guien matix is — A = (3%). The eigen values of A are given by - 1A-111=0 (3-1)3-1 = 0 » X-01 +8 =0 3) -21-41+8=0 = (- -+d=o > c=d. so, vz = (). Matrix P that diagonalizes A orthogonally is - P = (viva) 6 :). Novo, per poi (5) -=(1-33-6-3The eigen values of A are given by - 1A-1I[=0 3 16- 2V3 -0 12V3 71 =) (6-4) (7-1) - 12 = 0 > -131+42-12 =0 =) 2-131+ 30 =0 -1ELL.LU let V2 = (i) be an eigen rector corresponding to 12 = 10 Then, (A-101) V2 = 0 13. (2)()-(3) >(v 3)(5)=68) [RPGSM * 6 o